# FreDN: Spectral Disentanglement for Time Series Forecasting via Learnable Frequency Decomposition

### v1

@@ 14. $\hat{Y}=\operatorname{RevIN}^{-1}(\hat{Y}_{\text{season}}+\hat{Y}_{\text{trend}})$
Horizon ETTh1 ETTh2 ETTm1 ETTm2
MSE MAE MSE MAE MSE MAE MSE MAE
@@
Horizon Weather Electricity Traffic
MSE MAE MSE MAE MSE MAE
@@
Appendix Appendix F Visualization
W present supplementary prediction visualizations on all seven datasets in Figures 16, 17, 18, 19, 20, 21, 22, covering the following representative baselines: FreDF (wang2025fredf), SOFTS (han2024softs), TimeMixer (wang2024timemixer), PatchTST (nie2023patchtst), and DLinear (zeng2023transformers), alongside our proposed FreDN. To highlight the strength in leveraging long-range temporal information, we set the input sequence length to $L=720$. For layout clarity, we fix the prediction horizon to $\tau=96$ and visualize only the last $\tau$ time steps of the input. Among all methods, FreDN delivers the most coherent and accurate forecasts.
Appendix Appendix G Limitation and Future Work
@@
Appendix Appendix H Societal Impacts

### v2

@@ Appendix E Error Bar
Datasets ETTh1 ETTh2 ETTm1 ETTm2
Horizon MSE MAE MSE MAE MSE MAE MSE MAE
@@
Datasets Weather Electricity Traffic
Horizon MSE MAE MSE MAE MSE MAE
@@
Appendix F Visualization
W present supplementary prediction visualizations on all seven datasets in Figures 16, 17, 18, 19, 20, 21, 22, covering the following representative baselines: FreDF (Wang et al. 2025a), SOFTS (Han et al. 2024), TimeMixer (Wang et al. 2024), PatchTST (Nie et al. 2023), and DLinear (Zeng et al. 2023), alongside our proposed FreDN. To highlight the strength in leveraging long-range temporal information, we set the input sequence length to $L=720$. For layout clarity, we fix the prediction horizon to $\tau=96$ and visualize only the last $\tau$ time steps of the input. Among all methods, FreDN delivers the most coherent and accurate forecasts.
Appendix G Limitation and Future Work
@@
Appendix H Societal Impacts